When the price of DVDs falls from 10 to 8, the quantity supplied of DVDs falls from 398 to 216. What is the price elasticity of supply of DVDs? (Enter your answer as a decimal rounded to two decimal places, not a fraction)
Answer #1

Price elasticity of supply = percentage change in supply quantity/ percentage change in price

Percentage change in price = 8-10/10 = -2/10 = -20%

Percentage change in supply = 216-398/398 = -45.72%

Price elasticiy of supply = 20%/45.72% = 0.43744

Price elasticty of supply = 43.74%